OpenStudy (anonymous):

I have a couple of questions I need help solving them please don't give me the answer. -Solve 4x + 3y = 8x + y for y -Which of the following ordered pairs is a solution of y < 2x -6?

OpenStudy (anonymous):

Hello there, I think I can help you!

OpenStudy (anonymous):

Please .. I don't want the answers I just want help getting there.

OpenStudy (anonymous):

Sure thing, what do you seem to be having trouble with?

OpenStudy (anonymous):

Really starting the problems ...

OpenStudy (anonymous):

About those two problems.

OpenStudy (anonymous):

Yes ..

OpenStudy (anonymous):

Ah, okay, well let's look at the first one.

OpenStudy (anonymous):

Ok... I keep writing every way but, none of my ways match the answers ....

OpenStudy (anonymous):

4x + 3y = 8x + y for y The goal here is to get y all on one side, and get all the x's on the other.

OpenStudy (anonymous):

So what do I add the 4x+8x and the 3y+y ?

OpenStudy (anonymous):

\[ 4x + 3y = 8x + y\]\[ 4x + 3y - y = 8x\]\[ 4x + 2y = 8x\]\[2y = 8x - 4x\]\[2y = 4x\]\[y = 4x/2\]\[y = 2x\]

OpenStudy (anonymous):

I posted the steps one by one. Do you need a numerical value for y? or just one that contains x?

OpenStudy (anonymous):

Okay .. I see exactly what I did wrong ...

OpenStudy (anonymous):

Ah, glad to hear that.

OpenStudy (anonymous):

Now the next one?

OpenStudy (anonymous):

Yes me too .. Because I would've guessed ... :o We don't want that

OpenStudy (anonymous):

Which of the following ordered pairs is a solution of y < 2x -6?

OpenStudy (anonymous):

Can you list the ordered pairs?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Here you go (0,5) (2, -3) (5, 8)

OpenStudy (anonymous):

Okay, so the goal here is to plug in each ordered pair into y < 2x -6 and see which one makes that statement true.

OpenStudy (anonymous):

Ok... So how do we do that .. Do we solve the problem ?

OpenStudy (anonymous):

So, for the first one, we plug in 0 to x, and 5 to y, and get 5 < 2(0) - 6, which comes out to be 5 < -6, which is false because -6 is less than 5, so we throw that one out and go to the next one.

OpenStudy (anonymous):

Ohhhh ... I see.

OpenStudy (anonymous):

So it's b?

OpenStudy (anonymous):

y < 2x -6, -3 < 4 - 6, -3 < -2 which is true, so yeah looks like it.

OpenStudy (anonymous):

You are amazing ... Thank you so much !

OpenStudy (anonymous):

Thank you, and good luck. :)